OpenStudy (anonymous):

A 100 mg sample of compound CxHyOz (where x,y, and z are subscripts) was burned to form 199.8 mg CO2 and 81.8 mg water. Find the empirical formula

OpenStudy (xishem):

First, try writing out the chemical equation for the reaction.

OpenStudy (anonymous):

CxHyOz >> xCO2+ y/2H2O since the number of moles in both side of the equation must be equal x moles of C in CxHyOz will give x moles of CO2 y moles of H in CxHyOz will give y/2 moles of H2O (because there's 2 moles of H in one mole of H2O) so back to the calculations: m(CO2)=0.1998g m(H2O)=0.0818g n(CO2)=m/M n(H2O)=m/M = 0.1998/(12.01+2x16) =0.0818/(2x1.008+16) =4.540x10^-3 mol = 4.54x10^-3mol so n(C in CxHyOz ) = n(CO2) =4.540x10^-3 mol n(H in CxHyOz ) = 2n(H20) = 9.081x10^-3 mol n(O in CxHyOz)=2n(CO2) + n(H2O) =1.36x10^-2 mol for so ratio in C:H:O = 4.540x10^-3 : 9.081x10^-3 : 1.36x10^-2 = 1 : 2 : 3 the empirical formula for CxHyOz is CH2O3 (H2CO3, the substance is carbonic acid)